1.y=x^2+3
2.3-4x=y
Find domain and range of both one and two

Answer by stanbon(48535)

1. y=x^2+3
Domain: x can take any real number value.
Range: Since x^2 is always >=0 x^2+3 is
always >=3. So, range is >=3

2. 3-4x=y
Domain: x can take any real number value.
Range: y will cover all real number values.
Cheers,
Stan H.